A fairly priced unlevered firm plans to pay a dividend of $2 next year (t=1) which is expected to grow by 5% pa every year after that. The firm's required return on equity is 12% pa. The firm is thinking about reducing its future dividend payments by 10% so that it can use the extra cash to invest in more projects which are expected to return 12% p.a., and have the same risk as the existing projects. Therefore, next year's dividend will be $1.80. What will be the stock's new annual capital return (proportional increase in price per year) if the change in payout policy goes ahead? Assume that payout policy is irrelevant to firm value and that all rates are effective annual rates.
Answer #1

Current situation

D1 = $ 2; g = 5%; Ke = 12%

Hence, price of the stock, P = D1 / (Ke - g) = 2 / (12% - 5%) = $ 28.57

Revised situation:

New D1 = 2 x (1 - 10%) = $ 1.80

Since, we are assuming that payout policy is irrelevant to firm value

Price of the stock = 28.57 (same as before). Let the new growth rate be g1.

Hence, Price = 28.57 = New D1 / (Ke - g1) = 1.80 / (12% - g1)

Hence, g1 = 12% - 1.80 / 28.57 = 5.70%

Hence, the stock's new annual capital return (proportional increase in price per year) if the change in payout policy goes ahead = g1 = 5.70%